# Bright traveling breathers in media with long-range, nonconvex dispersion

### v2

@@ IV.2 Conduit equation bright traveling breather solutions
The examination of BBM bright traveling breathers has primed us for an investigation of their analogs in the conduit equation, which has an identical unit mean linear dispersion relation.
@@
The computations are initialized using the weakly nonlinear NLS approximation (16) with $\tilde{k}=4$ and $\tilde{a}\approx 0.38$. As before, the amplitude scale in Eq. (16) is defined by $\epsilon=(\tilde{a}\gamma)/4$, with $\gamma=(8\tilde{k}^{4}+5\tilde{k}^{2}+3)/(3\tilde{k}^{5}+12\tilde{k}^{3}+9\tilde{k})$. While recovering the branch of traveling breathers, we observe the familiar trends associated with a shifting wave mean and the emergence of a periodic carrier background. After this family, referred to as Family-1, is computed, we applied the scaling symmetry (10) to normalize all the breathers to unit mean. The continuation runs slow down once the number of cycles of carrier wave oscillations in the breather core limit to three, in contrast to the continuation runs for BBM breathers, where the limit was at one cycle. Like in the case of BBM breathers, the group velocity $\partial_{\tilde{k}}\omega_{0}=2(1-\tilde{k}^{2})/(1+\tilde{k}^{2})^{2}$ in the weakly nonlinear regime appears to be an upper limit for conduit breathers. This precludes breathers propagating with positive velocity. Moreover, a lower limit for breather velocity is not apparent from the computational results. On the other hand, a point of difference with the BBM waveforms is a pronounced asymmetry, which tends to bound the conduit wave profiles away from zero.
@@
To illustrate these trends, we present a few representative conduit breathers from Family-1 in Fig. 7. The relevant computations are performed on a spatial domain of length $1202L=$ 120 with spatial step $\Delta x=\frac{2L}{2M}\approx 0.06$ while $N=16$ ($32$ Fourier modes) provides the temporal discretization. The entire branch is contained within a relatively small velocity interval of width $\approx 0.012$.
@@
Figure 7: The Family-1 of traveling breathers of the conduit equation with $(T,c)\approx$ (a) $7.172,-0.101$, (b) $7.093,-0.106$, (c) $7.079,-0.107$, (d) $7.073,-0.109$, (e) $7.072,-0.111$, (f) $7.116,-0.113$.
@@
Figure 8: Variation of the (a) periodic background amplitude (23), (b) breather amplitude (24), (c) periodic background wavenumber, (d) phase shift $\sigma$ (26) and (e) time-period $T$ with breather velocity $c$ for Family-1 of conduit breathers.
@@
Figure 9: Computed unit-mean traveling breathers from the families (filled circles) and the weakly nonlinear prediction $T=\mathcal{T}$ (28) (colored surface).
@@
Figure 10: Family-2 of conduit weakly nonlinear wavepackets near the zero dispersion line, characterized by (a) carrier wave number $\tilde{k}\approx 2$ and $(\tilde{a},c,T)\approx 0.166,-0.240,4.907$, (b) $\tilde{k}\approx 1.8$ and $(\tilde{a},c,T)\approx 0.210,-0.243,4.886$.
@@
In Fig. 8, we show how the identifying parameters ${a}$, $\tilde{a}$, ${k}$, $\sigma$ and $T$ vary with the breather velocity $c$ for Family-1, upon the emergence of the periodic background. With an increasing magnitude of the wavepacket velocities, increasing trends in ${a}$ and $\tilde{a}$ are observed, while ${k}$ displays slight variation. The phase jump $\sigma$ is negative for all breathers in this branch. Given the strongly nonlinear nature of the conduit equation, it is interesting to check how close the traveling breathers are to the NLS regime. To this end, as before, the time-period $T$, velocity $c$ and amplitude $\tilde{a}$ form a set of identifying parameters. Moreover, the time-period in the NLS approximation is given in Eq. (28), where, for the conduit equation, the linear frequency $\Omega_{0}(c)$ is
@@
thus necessitating $-1/4<c_{0}<0$ for a real frequency. Additionally, $\gamma(c)$ is defined to be
@@
×$\times\frac{(16+16\sqrt{4c+1}+43c+11c\sqrt{4c+1}+6c^{2})}{(2+2\sqrt{4c+1}+2c-2c\sqrt{4c+1})}.$
The NLS surface for the conduit equation $\mathcal{T}(\tilde{a},c)$ is shown in Fig. 9. For appreciably small intervals in velocity, time-period and breather amplitudes, the wavepackets reside on the $\mathcal{T}(\tilde{a},c)$ surface, and thereafter lie entirely below this.
We also investigate the existence and form of weakly nonlinear conduit breathers near the zero-dispersion line. To this end, we initiate the computations with an appropriate NLS initial guess. We recover a delocalized wavepacket with a nearly monochromatic amplitude-modulated core with $\tilde{k}=2$ (see Fig. 10.a). We seed the $c$-continuation algorithm with this wavepacket, to obtain the breather in Fig. 10 (b). This wavepacket is characterized by core carrier wave oscillations with wavenumber $\tilde{k}=1.8$. The continuation procedure is seen to slow down significantly thereafter. It is remarkable that even for such small amplitude traveling breathers, there is a relatively large amplitude periodic background, which points to the operating higher-order dispersive effects therein. At this point, the existence of bright wavepackets across the zero-dispersion line is unclear but is interesting for future investigation.
@@ IV.3 Dynamic stability of breathers
We numerically investigate the dynamic stability of computed BBM and conduit traveling breathers with direct numerical simulations. The initial condition consists of a numerically computed traveling breather solution $1+\beta(\chi,\tau)$ evaluated at $\tau=0$ that is multiplicatively perturbed: $u(x,0)=(1+\beta(x,0))(1+\Delta(x))$ for the BBM equation and similarly for the conduit equation. A smooth perturbation function $\Delta$ is constructed from a spatially periodic disturbance with random Fourier series coefficients that is band-limited and scaled to have a peak amplitude of $0.05$. We employ the standard fourth-order Runge-Kutta explicit time-stepping scheme along with a Fourier discretization in space similar to Maiden and Hoefer (2016); Congy et al. (2020). We perform long-time numerical integration for more than 100 breather periods $T$ of two perturbed traveling breather solutions, one weakly nonlinear and the other strongly nonlinear, in each of the seven computed wave families (five BBM and two conduit families). All exhibit similar dynamical behavior. A representative example of a numerically evolved, perturbed strongly nonlinear breather solution from Family-1 of the conduit equation compared with the unperturbed breather is shown in Fig. 11 (a) and (b). The traveling breather core is slightly delayed after an evolution time of $150T$. Despite the 5% initial perturbation and long evolution time, the breather retains its coherence.
@@
Figure 11: Evolution of a perturbed strongly nonlinear conduit traveling breather from Family-1. Sub-figure (a) shows a contour plot describing the spatiotemporal structure of the perturbed waveform when evolved to $t=150T\approx 1067$. (b) Time snapshots of the perturbed (solid red) and unperturbed waveforms (dashed blue) are shown at the end time, pointing to the coherence of the breather.
V Discussions and Conclusions
@@
Branches of bright traveling breathers have been computed as solutions to a space-time boundary value problem for the BBM and conduit equations. For both, we found that traveling breathers are approximated by NLS envelope bright solitons for small amplitudes and carrier wavenumbers sufficiently deep in the negative dispersion regime. An emergent feature of traveling breathers in the weakly nonlinear regime was de-localization, signified by the presence of a propagating, periodic background. While the NLS approximation remains accurate in the traveling breather core, it is necessary to examine the effects of higher-order dispersive corrections to explain the de-localization, which we briefly describe now.
@@
A convenient framework to understand de-localization is the third-order NLS (TNLS):
[…]
@@
where $\epsilon^{\prime}={\epsilon}({\partial_{\tilde{k}\tilde{k}\tilde{k}}\omega_{0}})/({6(\partial_{\tilde{k}\tilde{k}}\omega_{0})^{3/2}})$, the slow, traveling spatial coordinate $\zeta$ is defined in Eq. (12) and $\epsilon$ is the amplitude scale (cf. eq. (11)). Without loss of generality, we consider the usual unit amplitude NLS envelope soliton $B(\zeta,\tau)={\rm sech}(\zeta)\exp(i\tau/2)$. Linearizing (31) about the soliton and seeking a resonant solution with the same frequency of the form $\exp\left(i(\kappa{\zeta}+\frac{{\tau}}{2})\right)$, we obtain the cubic equation for $\kappa$
@@
which admits exactly one short wave solution $\kappa_{0}\sim-\frac{1}{2\epsilon^{\prime}}$, $|\kappa_{0}|\gg 1$ as $\epsilon{\prime}\to 0$. This resonance with the linear spectrum has two implications: (a) NLS solitons radiate short waves when subject to weak third-order dispersion, (b) steady solutions to Eq. (31), $B({\zeta},{\tau})\equiv A\left({\zeta}-\epsilon^{\prime}c^{\prime}{\tau}\right)\exp{i({\tau}/2)}$ are de-localized. Both the unsteady and steady problems are studied under the framework of exponential asymptotics (see Yang (2010)), which yields estimates for the amplitudes of the one-sided short wave radiation emitted by NLS solitons and the far-field (${\zeta}\rightarrow\pm\infty$) tails of the steady, de-localized waveforms.
@@
An examination of the steady ODE problem reveals that $A(\xi)$ must be exponentially small in $\epsilon^{\prime}$ as $|\xi|\to\infty$. The complex-valued profile $A(\xi)$ exhibits a non-zero phase jump. The implications of this perturbative analysis for weakly nonlinear traveling breathers are that they generically admit a parameterization in terms of the carrier wave number, traveling breather amplitude, and phase jump, respectively. Moreover, this exponential asymptotic result corroborates our computational finding of small-amplitude traveling breathers with nearly localized, spatial waveforms.
@@
The wavenumber of the resonant wave in the fast spatial coordinate $x$ is $\tilde{k}_{1}\equiv\tilde{k}+\kappa_{0}\frac{\epsilon}{\sqrt{\partial_{\tilde{k}\tilde{k}}\omega_{0}}}\sim\tilde{k}-3(\partial_{\tilde{k}\tilde{k}}\omega_{0}/\partial_{\tilde{k}\tilde{k}\tilde{k}}\omega_{0})$. This TNLS prediction is viable provided $\partial_{\tilde{k}\tilde{k}\tilde{k}}\omega_{0}\neq 0$. We have compared this wavenumber prediction for the BBM and conduit dispersion with the computed traveling breather solutions and observed a significant discrepancy in the interval $\tilde{k}\in[3,3.5]$ for nearly monochromatic BBM (or conduit breathers). The relative errors here were found to be as large as $1600\%$. On the other hand, in the interval $\tilde{k}\in[1.74,1.9]$, the relative errors were contained below $20\%$ and were found to be as low as $6\%$ for $\tilde{k}\approx 1.8$. This discrepancy is attributed to the intricate structure of the dispersion relations (3), (9) for which $\partial_{\tilde{k}\tilde{k}\tilde{k}}\omega_{0}$ is zero when $\tilde{k}=\sqrt{2}+1$ while the dispersion curvature has a zero at $\tilde{k}=\sqrt{3}$. We suspect that to complete the characterization of the traveling breather periodic background, a cubic NLS model incorporating the full-dispersion of these non-local models may be required.
@@
Another implication of the BBM/conduit non-convex dispersion relation is the persistence of traveling bright breathers in the weakly nonlinear defocusing regime (cf. Fig. 5). A preliminary insight into this persistence can be gained within a TNLS framework Kivshar and Luther-Davies (1998); Kivshar and Afanasjev (1991), where it was shown that traveling breathers near the zero-dispersion point manifest a bright or anti-dark waveform instead.
@@
In summary, we have introduced a direct computational method for traveling bright breathers of nonlinear dispersive equations. Multiple families of BBM and conduit equation traveling breathers have been obtained. In the weakly nonlinear regime, these limit to amplitude-modulated wavepackets that are well-approximated by the NLS equation. In the strongly nonlinear regime, these traveling breathers were seen to be de-localized, bright modulation defects on cnoidal-type carrier waves. Large amplitude BBM breathers were seen to have more pronounced cnoidal backgrounds than the conduit counterparts we computed. Our computations indicate that BBM and conduit bright traveling breathers bifurcating from NLS bright solitons propagate with negative velocities only, and thus, it is required to turn to an alternate setup than what is currently being employed, for their experimental generation Mao et al. (2023). Finally, BBM and conduit traveling breather solutions were found to be dynamically stable over the course of long-time numerical evolution of their initially perturbed waveforms. A more detailed study investigating the stability of these traveling breathers through the lens of Floquet theory is possible future work. Another interesting problem is the experimental generation of bright breather trains and of even a breather gas, from an unstable periodic wave Maiden and Hoefer (2016). The latter could also form the basis for future investigations in other relevant geophysical Whitfield and Johnson (2015); Helfrich (2007) or short-pulse optical scenarios Leblond and Mihalache (2013); Costanzino et al. (2009)
@@
Yet another extension of the present work is the computation of bright and dark traveling breathers Hoefer et al. (2023) which are generated in the KdV equation via the interaction of solitons and cnoidal waves Kuznetsov and Mikhailov (1975). Such classes of traveling breathers have been observed experimentally over wide amplitude ranges in Mao et al. (2023), and have been seen to exhibit qualitatively similar properties as their asymptotic, KdV reductions. An open question is how these solutions relate to the bright breathers computed here. Can soliton-cnoidal wave interaction solutions be continued to the bright traveling breather solutions obtained here that bifurcate from bright soliton solutions of the focusing NLS equation? Our computational method could help establish the existence and properties of such dark and bright traveling breathers in the absence of integrable structure. A tantalizing problem is the existence of more general two-phase solutions that could also be explored using a similar computational framework.

### v3

@@ IV.2 Conduit equation bright traveling breather solutions
The examination of BBM bright traveling breathers has primed us for an investigation of their analogs in the conduit equation, which has an identical unit-mean linear dispersion relation.
@@
The computations are initialized using the weakly nonlinear NLS approximation (16) with $\tilde{k}=4$ and $\tilde{a}\approx 0.38$. As before, the amplitude scale in Eq. (16) is defined by $\epsilon=\tilde{a}\gamma/4$, with $\gamma=(8\tilde{k}^{4}+5\tilde{k}^{2}+3)/(3\tilde{k}^{5}+12\tilde{k}^{3}+9\tilde{k})$. While continuing the branch of traveling breathers, we observe the familiar trends associated with a shifting wave mean and the emergence of a periodic carrier background. After this family, referred to as family 1, is computed, we apply the scaling symmetry (10) to normalize all the breathers to unit mean. The continuation runs slow down once the number of cycles of carrier wave oscillations in the breather core limit to three, in contrast to the continuation runs for BBM breathers, where the limit was at one cycle. Like in the case of BBM breathers, the group velocity $\partial_{\tilde{k}}\omega_{0}=2(1-\tilde{k}^{2})/(1+\tilde{k}^{2})^{2}$ in the weakly nonlinear regime appears to be an upper limit for conduit breathers. This precludes breathers propagating with positive velocity because $\tilde{k}>\sqrt{3}>1$. Moreover, a lower limit for breather velocity is not apparent from the computational results. On the other hand, a point of difference with the BBM waveforms is a pronounced asymmetry, which tends to bound the conduit wave profiles away from zero.
@@
To illustrate these trends, we present a few representative conduit breathers from family 1 in Fig. 7. The relevant computations are performed on a spatial domain of length $1202L=$ 120 with spatial step $\Delta x=\frac{2L}{2M}\approx 0.06$ while $N=16$ ($32$ Fourier modes) provides the temporal discretization. The entire branch is contained within a relatively small velocity interval of width approximately equal to $0.012$.
@@
Figure 7: Family 1 of traveling breathers of the conduit equation with (a) $(T,c)\approx 7.172,-0.101$, (b) $(T,c)\approx 7.093,-0.106$, (c) $(T,c)\approx 7.079,-0.107$, (d) $(T,c)\approx 7.073,-0.109$, (e) $(T,c)\approx 7.072,-0.111$, (f) $(T,c)\approx 7.116,-0.113$.
@@
Figure 8: Variation of the (a) periodic background amplitude (23), (b) breather amplitude (24), (c) periodic background wave number, (d) phase shift $\sigma$ (26), and (e) timeperiod $T$ with breather velocity $c$ for family 1 of conduit breathers.
@@
Figure 9: Computed unit-mean traveling breathers from the families (closed circles) and the weakly nonlinear prediction $T=\mathcal{T}$ (28) (colored surface).
@@
Figure 10: Family 2 of conduit weakly nonlinear wavepackets near the zero-dispersion line, characterized by (a) carrier wave number $\tilde{k}\approx 2$ and $(\tilde{a},c,T)\approx 0.166,-0.240,4.907$ and (b) $\tilde{k}\approx 1.8$ and $(\tilde{a},c,T)\approx 0.210,-0.243,4.886$.
@@
In Fig. 8 we show how the identifying parameters ${a}$, $\tilde{a}$, ${k}$, $\sigma$, and $T$ vary with the breather velocity $c$ for family 1, upon the emergence of the periodic background. With an increasing magnitude of the wave-packet speed, increasing trends in ${a}$ and $\tilde{a}$ are observed, while ${k}$ displays slight variation. The phase jump $\sigma$ is negative for all breathers in this branch. Given the strongly nonlinear nature of the conduit equation, it is interesting to check how close the traveling breathers are to the NLS regime. To this end, as before, the timeperiod $T$, velocity $c$, and amplitude $\tilde{a}$ form a set of identifying parameters. Moreover, the timeperiod in the NLS approximation is given in Eq. (28), where, for the conduit equation, the linear frequency $\Omega_{0}(c)$ is
@@
thus necessitating $-\frac{1}{4}<c_{0}<0$ for a real frequency. Additionally, $\gamma(c)$ is defined to be
@@
×$\times\frac{16+16\sqrt{4c+1}+43c+11c\sqrt{4c+1}+6c^{2}}{2+2\sqrt{4c+1}+2c-2c\sqrt{4c+1}}.$
The NLS surface for the conduit equation $\mathcal{T}(\tilde{a},c)$ is shown in Fig. 9. For appreciably small intervals in velocity, timeperiod and breather amplitudes, the wave packets reside on the $\mathcal{T}(\tilde{a},c)$ surface and thereafter lie entirely below it.
We also investigate the existence and form of weakly nonlinear conduit breathers near the zero-dispersion line. To this end, we initiate the computations with an appropriate NLS initial guess. We recover a delocalized wavepacket with a nearly monochromatic amplitude-modulated core with $\tilde{k}=2$ [see Fig. 10.a]. We seed the $c$-continuation algorithm with this wavepacket, to obtain the breather in Fig. 10(b). This wavepacket is characterized by core carrier wave oscillations with wavenumber $\tilde{k}=1.8$. The continuation procedure is seen to slow down significantly thereafter. It is remarkable that even for such small-amplitude traveling breathers, there is a relatively large-amplitude periodic background, which points to the operable higher-order dispersive effects therein. At this point, the existence of bright wave packets across the zero-dispersion line is unclear but is interesting for future investigation.
@@ IV.3 Dynamic stability of breathers
We numerically investigate the dynamic stability of computed BBM and conduit traveling breathers with direct numerical simulations. The initial condition consists of a numerically computed traveling breather solution $1+\beta(\chi,\tau)$ evaluated at $\tau=0$ that is multiplicatively perturbed: $u(x,0)=[1+\beta(x,0)][1+\Delta(x)]$ for the BBM equation and similarly for the conduit equation. A smooth perturbation function $\Delta$ is constructed from a spatially periodic disturbance with random Fourier series coefficients that is band limited and scaled to have a peak amplitude of $0.05$. We employ the standard fourth-order Runge-Kutta explicit time-stepping scheme along with a Fourier discretization in space similar to Maiden and Hoefer (2016); Congy et al. (2020). We perform long-time numerical integration for more than 100 breather periods $T$ of two perturbed traveling breather solutions, one weakly nonlinear and the other strongly nonlinear, in each of the seven computed wave families (five BBM and two conduit families). All exhibit similar dynamical behavior. A representative example of a numerically evolved, perturbed strongly nonlinear breather solution from family 1 of the conduit equation compared with the unperturbed breather is shown in Figs. 11(a) and (b). The traveling breather core is slightly delayed after an evolution time of $150T$. Despite the 5% initial perturbation and long evolution time, the breather retains its coherence.
@@
Figure 11: Evolution of a perturbed strongly nonlinear conduit traveling breather from family 1. (a) Contour plot describing the spatiotemporal structure of the perturbed waveform when evolved to $t=150T\approx 1067$. (b) Time snapshots of the perturbed (red solid line) and unperturbed waveforms (blue dashed line) at the end time, pointing to the coherence of the breather.
V Discussion
@@
Branches of bright traveling breathers have been computed as solutions to a space-time boundary-value problem for the BBM and conduit equations. For both we found that traveling breathers are approximated by NLS envelope bright solitons for small amplitudes and carrier wave numbers sufficiently deep in the negative dispersion regime. An emergent feature of traveling breathers in the weakly nonlinear regime was delocalization, signified by the presence of a propagating periodic background. While the NLS approximation remains accurate in the traveling breather core, it is necessary to examine the effects of higher-order dispersive corrections to explain the delocalization, which we briefly describe now.
@@
A convenient framework to understand delocalization is the third-order NLS (TNLS) equation
$iB_{{\tau}}+\frac{B_{{\zeta}{\zeta}}}{2}+|B|^{2}B=i\epsilon^{\prime}B_{{\zeta}{\zeta}{\zeta}},0<\epsilon^{\prime}\ll 1,$ (31)
@@
where $\epsilon^{\prime}={\epsilon}({\partial_{\tilde{k}\tilde{k}\tilde{k}}\omega_{0}})/{6(\partial_{\tilde{k}\tilde{k}}\omega_{0})^{3/2}}$, the slow, traveling spatial coordinate $\zeta$ is defined in Eq. (12) and $\epsilon$ is the amplitude scale [cf. eq. (11)]. Without loss of generality, we consider the usual unit amplitude NLS envelope soliton $B(\zeta,\tau)={\rm sech}(\zeta)\exp(i\tau/2)$. Linearizing (31) about the soliton and seeking a resonant solution with the same frequency of the form $\exp\left[i(\kappa{\zeta}+\frac{{\tau}}{2})\right]$, we obtain the cubic equation for $\kappa$,
@@
which admits exactly one short-wave solution $\kappa_{0}\sim-\frac{1}{2\epsilon^{\prime}}$, $|\kappa_{0}|\gg 1$ as $\epsilon^{\prime}\to 0$. This resonance with the linear spectrum has two implications: (a) NLS solitons radiate short waves when subject to weak third-order dispersion and (b) steady solutions to Eq. (31), $B({\zeta},{\tau})\equiv A\left({\zeta}-\epsilon^{\prime}c^{\prime}{\tau}\right)\exp{i({\tau}/2)}$ are delocalized. Both the unsteady and steady problems are studied under the framework of exponential asymptotics (see Yang (2010)), which yields estimates for the amplitudes of the one-sided short-wave radiation emitted by NLS solitons and the far-field (${\zeta}\rightarrow\pm\infty$) tails of the steady, delocalized waveforms.
@@
An examination of the steady ODE problem reveals that $A(\xi)$ must be exponentially small in $\epsilon^{\prime}$ as $|\xi|\to\infty$. The complex-valued profile $A(\xi)$ exhibits a nonzero phase jump. The implications of this perturbative analysis for weakly nonlinear traveling breathers are that they generically admit a parametrization in terms of the carrier wave number, traveling breather amplitude, and phase jump, respectively. Moreover, this exponential asymptotic result corroborates our computational finding of small-amplitude traveling breathers with nearly localized spatial waveforms.
@@
The wavenumber of the resonant wave in the fast spatial coordinate $x$ is $\tilde{k}_{1}\equiv\tilde{k}+\kappa_{0}\frac{\epsilon}{\sqrt{\partial_{\tilde{k}\tilde{k}}\omega_{0}}}\sim\tilde{k}-3(\partial_{\tilde{k}\tilde{k}}\omega_{0}/\partial_{\tilde{k}\tilde{k}\tilde{k}}\omega_{0})$. This TNLS prediction is viable provided $|\partial_{\tilde{k}\tilde{k}}\omega_{0}/\partial_{\tilde{k}\tilde{k}\tilde{k}}\omega_{0}|\ll 1$. We have compared this wavenumber prediction for the BBM and conduit dispersion with the computed traveling breather solutions and observed a significant discrepancy in the interval $\tilde{k}\in[3,3.5]$ for nearly monochromatic BBM (or conduit breathers). The relative errors here were found to be as large as $1600\%$. On the other hand, in the interval $\tilde{k}\in[1.74,1.9]$, the relative errors were contained below $20\%$ and were found to be as low as $6\%$ for $\tilde{k}\approx 1.8$. This discrepancy is attributed to the intricate structure of the dispersion relations (3) and (9) for which $\partial_{\tilde{k}\tilde{k}\tilde{k}}\omega_{0}$ is zero when $\tilde{k}=\sqrt{2}+1\approx 2.41$ while $\partial_{\tilde{k}\tilde{k}}\omega_{0}$ is zero at $\tilde{k}=\sqrt{3}\approx 1.73$. We suspect that to complete the characterization of the traveling breather periodic background, a cubic NLS model incorporating the full dispersion of these nonlocal models may be required.
@@
Another implication of the BBM-conduit nonconvex dispersion relation is the persistence of traveling bright breathers in the weakly nonlinear defocusing regime (cf. Fig. 5). A preliminary insight into this persistence can be gained within a TNLS framework Kivshar and Luther-Davies (1998); Kivshar and Afanasjev (1991), where it was shown that traveling breathers near the zero-dispersion point manifest a bright or antidark waveform instead.
VI Conclusion
@@
In summary, we have introduced a direct computational method for traveling bright breathers of nonlinear dispersive equations. Multiple families of BBM and conduit equation traveling breathers have been obtained. In the weakly nonlinear regime, these limit to amplitude-modulated wave packets that are well approximated by the NLS equation. In the strongly nonlinear regime, these traveling breathers were seen to be delocalized, bright modulation defects on cnoidal-type carrier waves. Large-amplitude BBM breathers were seen to have more pronounced cnoidal backgrounds than the conduit counterparts we computed. Our computations indicate that BBM and conduit bright traveling breathers bifurcating from NLS bright solitons propagate with negative velocities only and thus it is required to turn to an alternate setup to what is currently being employed, for their experimental generation Mao et al. (2023). Finally, BBM and conduit traveling breather solutions were found to be dynamically stable over the course of long-time numerical evolution of their initially perturbed waveforms. A more detailed study investigating the stability of these traveling breathers using Floquet theory is possible future work. Another interesting problem is the experimental generation of bright breather trains and of even a breather gas, from an unstable periodic wave Maiden and Hoefer (2016). The latter could also form the basis for future investigations in other relevant geophysical Whitfield and Johnson (2015); Helfrich (2007) or short-pulse optical scenarios Leblond and Mihalache (2013); Costanzino et al. (2009)
@@
Yet another extension of the present work is the computation of bright and dark traveling breathers Hoefer et al. (2023) which are generated in the KdV equation via the interaction of solitons and cnoidal waves Kuznetsov and Mikhailov (1975). Such classes of traveling breathers have been observed experimentally over wide amplitude ranges in Mao et al. (2023) and have been seen to exhibit qualitatively similar properties to their asymptotic, KdV reductions. An open question is how these solutions relate to the bright breathers computed here. Can soliton-cnoidal wave interaction solutions be continued to the bright traveling breather solutions obtained here that bifurcate from bright soliton solutions of the focusing NLS equation? Our computational method could help establish the existence and properties of such dark and bright traveling breathers in the absence of integrable structure. A tantalizing problem is the existence of more general two-phase solutions that could also be explored using a similar computational framework.